Determine which lines, if any, are parallel or perpendicular.
Line a: y = 3x + 6, Line b: 3x + y = 6, Line c: 3y = 2x + 18

Answer by nerdybill(7384)

Rewriting each of the given equation into the "point-slope" form of
y = mx + b
.
Line a: y = 3x + 6
.
Line b: 3x + y = 6
y = -3x + 6
.
Line c: 3y = 2x + 18
y = (2/3)x + 6
.
Since NO two lines have the same slope -- no lines are parallel.
And, NO two lines have slopes that are "negative reciprocals"
.
Therefore, NONE of the lines are parallel OR perpendicular to each other.